OpenStudy (anonymous):

Find the indefinite integral: (2^-2x)dx I used two different methods and got two different answers, so I'm wondering which method is correct.. 1) integral of a^u = (1/lna)(a^u) + C plugged it in and got (1/ln2)(2^-2x) +C 2) u substitution; u=-2x, du=-2dx = (-1/2)(2^u) = -(2^-2x)/(2ln2) +C

zepdrix (zepdrix):

Looks like method 2 worked out correctly :o Notice that method 2 is actually a `u-sub` and then you simply did step 1 after that.

zepdrix (zepdrix):

But I would encourage you to try to get comfortable `skipping` the u-substitution on problems like this. If it's too difficult, then fineeee. But you'll get it eventually.

zepdrix (zepdrix):

Example:\[\Large\bf\sf (e^{2x})'\quad=\quad 2 e^{2x}\] When you differentiate, the coefficient on x tells you to `multiply` by an extra 2 because of the chain rule. \[\Large\bf\sf \int\limits e^{2x}\;dx\quad=\quad \frac{e^{2x}}{2}\] When you integrate, the coefficient on x tells you to `divide` by an extra 2 to compensate for the chain rule would normally produce from the other direction.

zepdrix (zepdrix):

So in your problem, (if you feel more comfortable doing a u-sub, then fineeeee) you end up dividing by -2 because of your exponent.

OpenStudy (anonymous):

Ohhhhhhhhhhhhhhhh okay that makes WAY more sense, thanks! @zepdrix

OpenStudy (anonymous):

Does that apply to all integral rules, including the trig ones?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

when u skip u-substitution, u just need to make sure that the thing u cooked up differentiates back to the integrand. skipping u-sub is called "advanced guessing" you need to get use to advanced guessing... for simpler integrals like these

ganeshie8 (ganeshie8):

\(\Large\bf\sf \int\limits e^{2x}\;dx\quad=\quad \frac{e^{2x}}{2} \)

ganeshie8 (ganeshie8):

differentiate the "right side thing", u should get back the integrand u started wid.

ganeshie8 (ganeshie8):

its a good idea to check/confirm when u do "advanced guessing"

ganeshie8 (ganeshie8):

btw, to answer ur original question, it doesnt work for everything. it works only in special cases like simple exponential functions : \(\large \int a^{cx} dx\) simple trig functions : \(\large \int \cos (ax + b) dx\) etc...

ganeshie8 (ganeshie8):

anyways it comes wid practice... we dont have to delve too much on this in the beginning :)

OpenStudy (anonymous):

Okay, thanks so much :) @ganeshie8

ganeshie8 (ganeshie8):

np :)